ocket Products manufactures three types of remote-control devices: Economy, Standard, and Deluxe. The company, which uses activity-based costing, has identified five activities (and related cost drivers). Each activity, its budgeted cost, and related cost driver is identified below.

Activity Cost Cost Driver
Material handling $ 225,000 Number of parts
Material insertion 2,475,000 Number of parts
Automated machinery 840,000 Machine hours
Finishing 170,000 Direct labor hours
Packaging 170,000 Orders shipped
Total $ 3,880,000


The following information pertains to the three product lines for next year:

Economy Standard Deluxe
Units to be produced 10,000 5,000 2,000
Orders to be shipped 1,000 500 200
Number of parts per unit 10 15 25
Machine hours per unit 1 3 5
Labor hours per unit 2 2 2


Assume that Rocket is using a volume-based costing system, and the preceding overhead costs are applied to all products on the basis of direct labor hours. The overhead cost that would be assigned to the Standard product line is closest to:

Multiple Choice

  • $456,471.

  • $646,471.

  • $961,176.

  • $1,141,176.

  • None of the answers is correct.
Answer #1

Total direct labor hours

= Economy (10,000*2) + Standard (5,000*2) + Deluxe (2,000*2) = 20,000 + 10,000 + 4,000 = 34,000 dlh

Overhead rate = Overhead cost/Total direct labor hours

= 3,880,000/34,000 = 114.18

Overhead cost allocated to standard product line

= (5,000*2) * 114.18

= 1,141,176 (nearest)

Option D